How do I set up the following word problem into a mathematical equation:
A customer walks into an electronics store and buys five MP3 players and eight sets of headphones, paying $840. A second customer buys three MP3 players and four sets of headphones, and pays $480. How much does an MP3 player cost? How much does a set of headphones cost?

Answer by Edwin McCravy(20086)

How do I set up the following word problem into a mathematical equation:
A customer walks into an electronics store and buys five MP3 players and eight sets of headphones, paying $840. A second customer buys three MP3 players and four sets of headphones, and pays $480. How much does an MP3 player cost? How much does a set of headphones cost?


Five times the cost of an MP3 player PLUS eight times
the cost of a set of headphones EQUALS $840.

                       5M + 8H = 840

Three times the cost of an MP3 player PLUS four times 
the cost of a set of headphones EQUALS $480.

                       3M + 4H = 480

So we have this system of equations:

                       5M + 8H = 840
                       3M + 4H = 480

Can you solve the system of equations by elimination or 
substitution? If not post again asking how.

Answer M = $120,  H = $30

Edwin